# Large Language Models:

### v1

@@ C.2.2 Linear Regression with Large Language Model Labels as Covariates
[…]
@@
[…]
for $\Omega^{validation}=\sigma^{4}_{-V}\frac{1-\rho_{v}}{\rho_{v}}\sigma_{VW}^{2}$. Consequently, we can compare the limiting variances, and again observe that the bias-corrected regression coefficient has a smaller limiting variance if $\frac{1-\rho_{p}}{\rho_{p}}\sigma^{2}_{\widehat{V}W}+2\sigma_{\widehat{V}W}\leq\frac{1-\rho_{v}}{\rho_{v}}\left(\sigma_{VW}^{2}-\sigma_{\Delta W}^{2}\right)$. This can be satisfied provided the LLM outputs are sufficiently accurate for the gold-standard measurement.
@@ Appendix D Additional Monte Carlo Simulations based on Congressional Legislation
In this section, we report additional Monte Carlo simulations based on the data from the Congressional Bills Project (CBP). We first illustrate how the performance of the bias-corrected regression coefficient varies with the size of the validation data. We further illustrate that the performance of the bias-corrected regression when the economic concept is used as a covariate in the linear regression, as described in Appendix C.2.2.
@@ D.1 Varying the Size of the Validation Data
In Section 5.4 of the main text, we evaluated the performance of the plug-in regression coefficient against the bias-corrected estimator using a 10% validation sample. We explore how the performance of the bias-corrected regression coefficient varies as we vary the size of the validation sample.
As in the main text, for a given bill topic $V_{r}$, covariate $W_{r}$, and pair of large language model and prompting strategy, we randomly draw a sample of $5,000$ bills. On this random sample, we first calculate the plug-in regression coefficient $\widehat{\beta}$. We next randomly reveal the ground-truth label $V_{r}$ on 5%, 10%, 25%, and 50% of the random sample of $5,000$ bills, producing validation samples of varying sizes. We then calculate the bias-corrected coefficient $\widehat{\beta}^{debiased}$ on each validation sample. We repeat these steps for $1,000$ randomly sampled datasets, and we calculate the average bias of these alternative estimates for the target regression $\beta^{*}$ of the ground-truth concept $V_{r}$ on the chosen covariate $W_{r}$ on all 10,000 bills as well as the coverage of conventional confidence intervals. We repeat this exercise for each possible combination of bill topic $V_{r}$, linked covariate $W_{r}$, large language model $m$ (either GPT-3.5-Turbo or GPT-4o), and prompting strategy $p$. This allows us to summarize how the plug-in regression performs against the bias-corrected regression across a wide variety of possible regression specifications, choices of large language model and prompting strategies.
Figure A5 illustrates the distribution of normalized bias across possible combinations of bill topic $V_{r}$, linked covariate $W_{r}$, large language model $m$, and prompting strategy $p$, as the size of the validation sample changes. The top panels of Table 4(b) and Table 5(b) report summary statistics for labels produced by GPT-3.5-Turbo and GPT-4o respectively. While we often see severe biases for the plug-in regression, by contrast the bias-corrected regression coefficient is on average equal to the target regression coefficient for all sizes of the validation sample.
The bottom panels of Table 4(b) and Table 5(b) provide summary statistics of the coverage of conventional confidence intervals for the target regression. We see substantial coverage distortions for the plug-in regression, whereas the bias-corrected regression delivers approximately correct coverage for all sizes of the validation sample.
Finally, Figure A6 compares the mean square error of the bias-corrected coefficient versus the validation-sample only estimate of the target regression as we vary the size of the validation sample. The bias-corrected coefficient obtains noticeable improvements in mean square error for the validation proportions equal to 5% and 10%. The bias-corrected coefficient performs similarly to the validation-sample only estimator for validation proportions equal to 25% and 50%.
@@
[…]
Notes: The normalized bias reports the average bias of the plug-in regression coefficient $\widehat{\beta}$ and the bias-corrected coefficient $\widehat{\beta}^{debiased}$ for the target regression coefficient divided by their respective standard deviations across simulations. For each combination of model topic $V_{r}$, covariate $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the plug-in regression coefficient $\widehat{\beta}$ and the bias-corrected regression coefficient $\widehat{\beta}^{debiased}$. We vary the size of the validation sample over 5%, 10%, 25% and 50%. Results are averaged over $1,000$ simulations. We summarize the distribution of normalized bias and coverage across regression specifications, choice of large language model and prompting strategies. See Appendix Section D.1 for discussion.
@@
[…]
Notes: The normalized bias reports the average bias of the plug-in regression coefficient $\widehat{\beta}$ and the bias-corrected coefficient $\widehat{\beta}^{debiased}$ for the target regression coefficient divided by their respective standard deviations across simulations. The coverage reports the fraction of simulations in which a 95% nominal confidence interval centered around the plug-in regression coefficient $\widehat{\beta}$ and the bias-corrected coefficient $\widehat{\beta}^{debiased}$ cover the target regression coefficient $\beta^{*}$. For each combination of model topic $V_{r}$, covariate $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the plug-in regression coefficient $\widehat{\beta}$ and the bias-corrected regression coefficient $\widehat{\beta}^{debiased}$. We vary the size of the validation sample over 5%, 10%, 25% and 50%. Results are averaged over $1,000$ simulations. We summarize the distribution of normalized bias and coverage across regression specifications, choice of large language model and prompting strategies. See Appendix Section D.1 for discussion.
@@
Table A5: Summary statistics for normalized bias and coverage for Monte Carlo simulations on congressional legislation, varying the size of the validation sample (GPT-4o)
[…]
Notes: The normalized bias reports the average bias of the plug-in regression coefficient $\widehat{\beta}$ and the debiased coefficient $\widehat{\beta}^{debiased}$ for the target regression coefficient divided by their respective standard deviations across simulations. The coverage reports the fraction of simulations in which a 95% nominal confidence interval centered around the plug-in regression coefficient $\widehat{\beta}$ and the debiased coefficient $\widehat{\beta}^{debiased}$ cover the target regression coefficient. For each combination of model topic $V_{r}$, covariate $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the plug-in regression $\widehat{V}_{r}^{m,p}=\alpha+\beta W_{r}$ and the debiased regression coefficient. We vary the size of the validation sample over 5%, 10%, 25% and 50%. Results are averaged over $1,000$ simulations. See Appendix Section D.1 for discussion.
@@
[…]
Notes: For each combination of model topic $V_{r}$, covariate $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the bias-corrected regression coefficient $\widehat{\beta}^{debiased}$ and the validation-sample only regression coefficient $\widehat{\beta}^{*}$. We calculate the mean square error of $\widehat{\beta}^{debiased}$ and $\widehat{\beta}^{*}$ for the target regression. We vary the size of the validation sample over 5%, 10%, 25% and 50%, and we average the results over $1,000$ simulations. We summarize the distribution of average mean square error across regression specifications, choice of large language model and prompting strategies. See Appendix Section D.1 for discussion.
@@ D.2 Large Language Model Labels as Covariates
As a first step, we use the sample random sample of $10,000$ Congressional bills from the main text, and we now regress alternative linked economic variables on dummy indicators for the large language model’s labeled economic concept – in this case, the policy topic of the bill. For alternative dependent variables such as whether the bill’s sponsor was a Democrat, whether the bill originated in the Senate, and the DW1 score of the bill’s sponsor, we run the regression $W_{r}=\widehat{V}_{r}^{m,p}\beta+\epsilon$ for each possible pair of large language model $m$ and prompting strategy $p$. In Figure A7, each row considers a different regression for a linked covariate $W_{r}$ as the dependent variable, and each column plots the t-statistic for different large language model labels $\widehat{V}_{r}^{m,p}$ associated with alternative bill topics (e.g., Health, Banking, Finance and Domestic Commerce, Defense, Government Operations, and Public Lands and Water Management). For every combination of the linked variable $W_{r}$ and policy topic area, we see substantial variation in the t-statistics across alternative large language models and prompting strategies. Table A6 summarizes the coefficient estimates across models and prompts for each choice of labeled policy topic and the covariate.
@@
[…]
Notes: On 10,000 Congressional bills, we prompt GPT-3.5-Turbo and GPT-4o to label each description for its policy topic area using alternative prompting strategies. For each model $m$ and prompt $p$, we regress a linked variable $W_{r}$ (whether the bill’s sponsor was a Democrat, whether the bill originated in the Senate, and the DW1 score of the bill’s sponsor) on indicators $\widehat{V}_{r}^{m,p}$ for the large language model’s labeled policy topic (i.e., Health; Banking, Finance, and Domestic Commerce; Defense; Government Operations; Public Lands and Water Management; Other). In each subplot, the t-statistic estimates are sorted in ascending order for clarity. See Appendix Section D.2 for discussion.
@@
Democrat Health 0.636 0.642 0.614 0.651 0.604
Democrat Banking, Finance & Domestic Com. 0.582 0.582 0.567 0.597 0.604
Democrat Defense 0.586 0.586 0.576 0.598 0.604
Democrat Government Operations 0.601 0.598 0.588 0.620 0.604
Democrat Public Lands & Water Management 0.588 0.587 0.579 0.599 0.604
Senate Health 0.334 0.331 0.319 0.364 0.317
Senate Banking, Finance & Domestic Com. 0.304 0.305 0.293 0.314 0.317
Senate Defense 0.295 0.295 0.280 0.308 0.317
Senate Government Operations 0.290 0.291 0.280 0.301 0.317
Senate Public Lands & Water Management 0.391 0.390 0.383 0.404 0.317
DW1 Health -0.091 -0.095 -0.102 -0.076 -0.062
DW1 Banking, Finance & Domestic Com. -0.043 -0.043 -0.056 -0.027 -0.062
DW1 Defense -0.015 -0.016 -0.023 -0.009 -0.062
DW1 Government Operations -0.047 -0.048 -0.066 -0.033 -0.062
DW1 Public Lands & Water Management -0.024 -0.024 -0.033 -0.009 -0.062
@@
[…]
Notes: On 10,000 Congressional bills, we prompt GPT-3.5-Turbo and GPT-4o to label each description for its policy topic area using alternative prompting strategies. For each model $m$ and prompt $p$, we regress a linked variable $W_{r}$ (whether the bill’s sponsor was a Democrat, whether the bill originated in the Senate, and the DW1 score of the bill’s sponsor) on indicators for whether the large language model labeled a particular policy topic $1\{\widehat{V}_{r}^{m,p}=v\}$, focusing on Health, Banking, Finance & Domestic Commerce (“Banking”), Defense, Government Operations, and Public Lands & Water Management (“Public Lands”) The final column (“Sample Average”) reports the average of the linked variable $W_{r}$ across all Congressional bills. See Appendix Section D.2 for discussion.
We next explore whether these biases can be addressed by collecting a small validation sample and implementing the bias-corrected procedure described in Appendix C.2.2 can address these issues. We leverage the same Monte Carlo simulation design as described in Section 5.4.2 of the main text. For each linked variable $W_{r}$ and pair of large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ bills from our dataset of $10,000$ bills. On this random sample of $5,000$ bills, we first calculate the plug-in coefficients $\widehat{\beta}$ from regressing $W_{r}$ on $\widehat{V}_{r}^{m,p}$ (, for $\widehat{V}_{r}^{m,p}$ a vector of indicators for the labeled policy topic). We next randomly reveal the ground-truth label $V_{r}$ on 10\$ of our random sample of $5,000$ bills, which produces a validation sample. We calculate the bias-corrected coefficients $\widehat{\beta}^{debiased}$ as described in Appendix C.2.2. We repeat these steps for $1,000$ randomly sampled datasets. We repeat this exercise for each possible combination of linked variable $W_{r}$, large language model $m$ (either GPT-3.5-turbo or GPT-4o) and prompting strategy $p$. This allows to summarize how the plug-in regression performs against the bias-corrected regression across a wide variety of possible regression specifications, choices of large language model and prompting strategies.
Figure A8 and Table 7(b) summarizes our results. The plug-in regression suffers from substantial biases for almost all combinations of linked variable $W_{r}$, large language model $m$ (either GPT-3.5-turbo and GPT-4o) and prompting strategy $p$. By contrast, using the validation sample for bias correction effectively eliminates these biases. Furthermore, the bottom panel of Table 7(b) further illustrates the coverage comparison between the plug-in regression and the bias-corrected estimator — while confidence intervals centered at the plug-in regression are significantly distorted, bias-correction restores nominal coverage.
Finally, Figure A9 compares the mean square error of the bias-corrected regression against directly estimating the target regression on the validation sample. For many regression specifications, choices of language model and prompting strategies, we again find that the MSE of the bias-corrected regression is smaller than that of the validation-sample only regression.
@@
[…]
Notes: The normalized bias reports the average bias of the plug-in regression coefficient $\widehat{\beta}$ and the bias-corrected coefficient $\widehat{\beta}^{debiased}$ for the target regression coefficient divided by their respective standard deviations across simulations. For each combination of left hand side variable $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the plug-in regression $\widehat{\beta}$ and bias-corrected coefficient $\widehat{\beta}^{*}$ with the policy topic as a covariate using a 10% validation sample. Results are averaged over $1,000$ simulations. See Appendix Section D.2 for discussion.
@@
[…]
Notes: The normalized bias reports the average bias of the plug-in regression coefficient $\widehat{\beta}$ and the debiased coefficient $\widehat{\beta}^{debiased}$ for the target regression coefficient divided by their respective standard deviations across simulations. The coverage reports the fraction of simulations in which a 95% nominal confidence interval centered around the plug-in regression coefficient $\widehat{\beta}$ and the debiased coefficient $\widehat{\beta}^{debiased}$ cover the target regression coefficient. For each combination of left hand side variable $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the plug-in regression $\widehat{\beta}$ and bias-corrected coefficient $\widehat{\beta}^{*}$ with the policy topic as a covariate using a 10% validation sample. Results are averaged over $1,000$ simulations. See Appendix Section D.2 for discussion.
@@
[…]
Notes: For each combination of left hand side variable $W_{r}$, large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ Congressional bills and calculate the plug-in regression $\widehat{\beta}$ and bias-corrected coefficient $\widehat{\beta}^{*}$ with the policy topic as a covariate using a 10% validation sample. We calculate the mean square error of $\widehat{\beta}^{debiased}$ and $\widehat{\beta}^{*}$ for the target regression $\beta^{*}$. Results are averaged over $1,000$ simulations. We summarize the distribution of average mean square error across regression specifications, choice of large language model and prompting strategies. See Appendix Section D.2 for discussion.

### v3

@@ C.2.2 Linear Regression with Large Language Model Labels as Covariates
$\sqrt{N_{R}}\left(\widehat{\alpha}^{debiased}-\alpha^{*}\right)\xrightarrow{d}N(0,\Omega^{debiased})$
@@
$\sqrt{N_{R}}\left(\hat{\alpha}^{validation}-\alpha^{*}\right)\xrightarrow{d}N(0,\Omega^{validation})$
for $\Omega^{validation}=\sigma^{4}_{-V}\frac{1-\rho_{v}}{\rho_{v}}\sigma_{VW}^{2}$. Consequently, we can compare the limiting variances, and again observe that the bias-corrected regression coefficient has a smaller limiting variance if $\frac{1-\rho_{p}}{\rho_{p}}\sigma^{2}_{\widehat{V}W}+2\sigma_{\widehat{V}W}\leq\frac{1-\rho_{v}}{\rho_{v}}\left(\sigma_{VW}^{2}-\sigma_{\Delta W}^{2}\right)$. This can be satisfied provided the LLM’s errors in reproducing the existing measurement are sufficiently small.
@@ Appendix D Additional Monte Carlo Simulations based on Congressional Legislation
In this section, we report additional Monte Carlo simulations based on the data from the Congressional Bills Project (CBP; CAP). We first illustrate how the performance of the bias-corrected regression coefficient varies with the size of the validation data. We further illustrate that the performance of the bias-corrected regression when the economic concept is used as a covariate in the linear regression, as described in Appendix C.2.2.
@@ D.1 Varying the Size of the Validation Data
In Section 4.3 of the main text, we evaluated the performance of the plug-in regression coefficient against the bias-corrected estimator using a 5% validation sample. We explore how performance varies as we vary the size of the validation sample.
For a given bill topic $V_{r}$, covariate $W_{r}$, and pair of large language model and prompting strategy, we randomly draw a sample of $5,000$ bills. On this random sample, we first calculate the plug-in regression coefficient $\widehat{\beta}$. We next randomly reveal the ground-truth label $V_{r}$ on 2.5% (125 bills), 5% (250 bills), 10% (500 bills), 25% (1250 bills), and 50% (2500 bills) of the random sample of $5,000$ bills. We then calculate the bias-corrected coefficient $\widehat{\beta}^{debiased}$ on each validation sample. We repeat these steps for $1,000$ randomly sampled datasets, and we calculate the average bias of these alternative estimates for the target regression $\beta^{*}$ of the ground-truth concept $V_{r}$ on the chosen covariate $W_{r}$ on all 10,000 bills as well as the coverage of conventional confidence intervals. We repeat this exercise for each possible combination of bill topic $V_{r}$, linked covariate $W_{r}$, large language model $m$, and prompting strategy $p$. This allows us to summarize how the plug-in regression performs against the bias-corrected regression across a wide variety of possible regression specifications, choices of large language model and prompting strategies.
Appendix Figure A6 illustrates the distribution of normalized bias across possible combinations of bill topic $V_{r}$, linked covariate $W_{r}$, large language model $m$, and prompting strategy $p$, as the size of the validation sample changes. The top panels of Appendix Tables A4-A7 report summary statistics for labels produced by each model respectively. While we often see severe biases for the plug-in regression, by contrast the bias-corrected regression coefficient is on average equal to the target regression coefficient for all sizes of the validation sample.
The bottom panels of Appendix Tables A4-A7 provide summary statistics of the coverage of conventional confidence intervals for the target regression. We see substantial coverage distortions for the plug-in regression, whereas the bias-corrected regression delivers approximately correct coverage for all sizes of the validation sample.
Finally, Appendix Figure A7 compares the mean square error of the bias-corrected coefficient versus the validation-sample only estimate of the target regression as we vary the size of the validation sample. The bias-corrected coefficient obtains noticeable improvements in mean square error for the validation proportions equal to 2.5%, 5% and 10%. The bias-corrected coefficient performs similarly to the validation-sample only estimator for validation proportions equal to 25%, although it is likely unrealistic that the researcher would collect such large validation samples in an empirical application.
@@
Table A5: Summary statistics for normalized bias and coverage for Monte Carlo simulations on congressional legislation for GPT-4o, varying the size of the validation sample.
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.062 -1.612 1.566
5% 0.030 -1.646 1.567
10% 0.058 -1.624 1.545
25% 0.040 -1.569 1.533
50% 0.040 -1.619 1.572
Coverage
2.5% 0.907 0.570 0.954
5% 0.906 0.589 0.953
10% 0.909 0.558 0.954
25% 0.900 0.585 0.955
50% 0.905 0.583 0.954
(a) Plug-in regression
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.000 -0.041 0.051
5% -0.002 -0.052 0.054
10% -0.002 -0.066 0.049
25% -0.003 -0.050 0.055
50% 0.003 -0.048 0.047
Coverage
2.5% 0.895 0.843 0.925
5% 0.927 0.903 0.945
10% 0.939 0.925 0.952
25% 0.947 0.936 0.958
50% 0.948 0.934 0.957
(b) Debiased regression
Table A6: Summary statistics for normalized bias and coverage for Monte Carlo simulations on congressional legislation for GPT-5-mini, varying the size of the validation sample.
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.151 -2.072 2.113
5% 0.168 -2.079 2.092
10% 0.141 -2.017 2.116
25% 0.160 -2.058 2.085
50% 0.171 -2.021 2.123
Coverage
2.5% 0.771 0.379 0.952
5% 0.779 0.387 0.953
10% 0.781 0.361 0.951
25% 0.780 0.390 0.951
50% 0.785 0.377 0.956
(a) Plug-in regression
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.000 -0.050 0.042
5% 0.005 -0.052 0.060
10% -0.001 -0.052 0.054
25% -0.002 -0.058 0.062
50% 0.002 -0.052 0.054
Coverage
2.5% 0.902 0.850 0.929
5% 0.930 0.906 0.946
10% 0.942 0.927 0.955
25% 0.947 0.935 0.959
50% 0.946 0.935 0.958
(b) Debiased regression
Table A7: Summary statistics for normalized bias and coverage for Monte Carlo simulations on congressional legislation for GPT-5-nano, varying the size of the validation sample.
@@ D.2 Large Language Model Labels as Covariates
We use the same random sample of $10,000$ Congressional bills from the main text, and we now regress alternative linked economic variables on dummy indicators for the large language model’s labeled economic concept – in this case, the policy topic of the bill. For alternative dependent variables such as whether the bill’s sponsor was a Democrat, whether the bill originated in the Senate, and the DW1 score of the bill’s sponsor, we run the regression $W_{r}=\widehat{V}_{r}^{m,p}\beta+\epsilon$ for each possible pair of large language model $m$ and prompting strategy $p$. In Appendix Figure A8, each row considers a different regression for a linked covariate $W_{r}$ as the dependent variable, and each column plots the t-statistic for different large language model labels $\widehat{V}_{r}^{m,p}$ associated with alternative bill topics. For every combination of the linked variable $W_{r}$ and policy topic area, we see substantial variation in the t-statistics across alternative large language models and prompting strategies. Appendix Table A8 summarizes the coefficient estimates across models and prompts for each choice of labeled policy topic and the covariate.
@@
DW1 Health -0.096 -0.097 -0.105 -0.078 -0.062
DW1 Banking, Finance & Domestic Com. -0.043 -0.043 -0.053 -0.029 -0.062
DW1 Defense -0.016 -0.016 -0.025 -0.010 -0.062
DW1 Government Operations -0.043 -0.046 -0.064 -0.024 -0.062
DW1 Public Lands & Water Management -0.025 -0.026 -0.033 -0.013 -0.062
Democrat Health 0.643 0.646 0.617 0.655 0.604
Democrat Banking, Finance & Domestic Com. 0.579 0.579 0.568 0.593 0.604
Democrat Defense 0.588 0.588 0.577 0.602 0.604
Democrat Government Operations 0.594 0.596 0.571 0.614 0.604
Democrat Public Lands & Water Management 0.589 0.588 0.581 0.598 0.604
Senate Health 0.331 0.327 0.320 0.359 0.317
Senate Banking, Finance & Domestic Com. 0.299 0.298 0.282 0.313 0.317
Senate Defense 0.293 0.294 0.276 0.308 0.317
Senate Government Operations 0.292 0.292 0.282 0.305 0.317
Senate Public Lands & Water Management 0.386 0.386 0.372 0.398 0.317
@@
We next explore whether these biases can be addressed by collecting a small validation sample and implementing the bias-corrected procedure described in Appendix C.2.2 can address these issues. We leverage the same Monte Carlo simulation design as described in Section 4.3 of the main text.
For each linked variable $W_{r}$ and pair of large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ bills from our dataset of $10,000$ bills. On this random sample of $5,000$ bills, we calculate the plug-in coefficients $\widehat{\beta}$ byregressing $W_{r}$ on $\widehat{V}_{r}^{m,p}$ (for $\widehat{V}_{r}^{m,p}$ a vector of indicators for the labeled policy topic). We next randomly reveal the ground-truth label $V_{r}$ on 5% of our random sample of $5,000$ bills, which produces a validation sample. We calculate the bias-corrected coefficients $\widehat{\beta}^{debiased}$ as described in Appendix C.2.2. We repeat these steps for $1,000$ randomly sampled datasets. We repeat this exercise for each possible combination of linked variable $W_{r}$, large language model $m$ (either GPT-3.5-turbo or GPT-4o) and prompting strategy $p$. This allows to summarize how the plug-in regression performs against the bias-corrected regression across a wide variety of possible regression specifications, choices of large language model and prompting strategies.
Appendix Figure A9 and Appendix Table A9 summarizes our results. The plug-in regression suffers from substantial biases for almost all combinations of linked variable $W_{r}$, large language model $m$, and prompting strategy $p$. By contrast, using the validation sample for bias correction effectively eliminates these biases. Furthermore, the bottom panel of Appendix Table A9 further illustrates the coverage comparison between the plug-in regression and the bias-corrected estimator — while confidence intervals centered at the plug-in regression are significantly distorted, bias-correction restores nominal coverage.
Finally, Appendix Figure A10 compares the mean square error of the bias-corrected regression against directly estimating the target regression on the validation sample. For many regression specifications, choices of language model and prompting strategies, we again find that the MSE of the bias-corrected regression is smaller than that of the validation-sample only regression.
@@
Median 5% 95%
Normalized Bias
Plug-In 0.034 -1.615 1.828
Debiased 0.003 -0.058 0.058
Coverage
Plug-In 0.927 0.521 0.953
Debiased 0.932 0.906 0.955
(c) GPT-5-mini
Median 5% 95%
Normalized Bias
Plug-In 0.126 -1.664 2.172
Debiased -0.002 -0.064 0.058
Coverage
Plug-In 0.900 0.405 0.950
Debiased 0.934 0.908 0.958
(d) GPT-5-nano
@@
D.2.1 Varying the Size of the Validation Data:
We evaluated the performance of bias-correcting linear regression that use large language model labels as covariates using a 5% validation sample. We finally explore how the performance of the bias-corrected regression coefficient varies as we vary the size of the validation sample. We repeat our Monte Carlo simulations now varying the size of the validation sample by randomly revealing the measurements $V_{r}$ on 2.5% (125 bills), 5% (250 bills), 10% (500 bills), 25% (1250 bills), and 50% (2500 bills) of the random sample of $5,000$ bills. The results are summarized in Appendix Figure A11, Appendix Tables A10-A13 and Appendix Figure A12. We continue to find that the bias-corrected regression performs well in finite samples, even when the validation sample only contains $125$ bills.
Figure A11: Normalized bias of the plug-in regression and bias-corrected regression using policy topic as a covariate as the validation sample size varies.
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.120 -1.376 2.030
5% 0.144 -1.514 2.083
10% 0.147 -1.437 2.099
25% 0.136 -1.469 2.024
50% 0.127 -1.486 2.043
Coverage
2.5% 0.900 0.373 0.951
5% 0.901 0.363 0.950
10% 0.897 0.353 0.949
25% 0.893 0.391 0.948
50% 0.893 0.368 0.951
(a) Plug-in regression
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.000 -0.057 0.050
5% 0.003 -0.051 0.060
10% -0.002 -0.055 0.057
25% -0.001 -0.053 0.052
50% -0.002 -0.055 0.056
Coverage
2.5% 0.909 0.868 0.959
5% 0.933 0.907 0.955
10% 0.941 0.927 0.954
25% 0.946 0.935 0.958
50% 0.947 0.936 0.959
(b) Debiased regression
Table A10: Summary statistics for normalized bias and coverage using policy topic as a covariate for GPT-3.5-Turbo, varying the size of the validation sample.
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.038 -1.087 1.509
5% 0.042 -1.146 1.558
10% 0.024 -1.118 1.624
25% 0.039 -1.145 1.541
50% 0.029 -1.111 1.490
Coverage
2.5% 0.928 0.675 0.954
5% 0.928 0.640 0.957
10% 0.927 0.642 0.952
25% 0.926 0.639 0.953
50% 0.922 0.648 0.953
(a) Plug-in regression
Validation Prop. Median 5% 95%
Normalized Bias
2.5% -0.003 -0.065 0.052
5% -0.003 -0.063 0.053
10% -0.001 -0.059 0.055
25% -0.002 -0.062 0.053
50% -0.003 -0.057 0.052
Coverage
2.5% 0.904 0.860 0.948
5% 0.930 0.900 0.952
10% 0.943 0.928 0.952
25% 0.947 0.933 0.957
50% 0.949 0.935 0.959
(b) Debiased regression
Table A11: Summary statistics for normalized bias and coverage using policy topic as a covariate for GPT-4o, varying the size of the validation sample.
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.031 -1.615 1.884
5% 0.034 -1.615 1.828
10% 0.035 -1.604 1.901
25% 0.041 -1.622 1.857
50% 0.039 -1.596 1.845
Coverage
2.5% 0.926 0.491 0.955
5% 0.927 0.521 0.953
10% 0.928 0.510 0.955
25% 0.930 0.505 0.953
50% 0.929 0.494 0.956
(a) Plug-in regression
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.005 -0.049 0.065
5% 0.003 -0.058 0.058
10% -0.004 -0.052 0.059
25% -0.001 -0.050 0.052
50% 0.003 -0.054 0.055
Coverage
2.5% 0.904 0.852 0.958
5% 0.932 0.906 0.955
10% 0.940 0.926 0.953
25% 0.947 0.935 0.958
50% 0.947 0.935 0.958
(b) Debiased regression
Table A12: Summary statistics for normalized bias and coverage using policy topic as a covariate for GPT-5-mini, varying the size of the validation sample.
Validation Prop. Median 5% 95%
Normalized Bias
2.5% 0.118 -1.631 2.112
5% 0.126 -1.664 2.172
10% 0.119 -1.635 2.201
25% 0.133 -1.695 2.182
50% 0.158 -1.638 2.177
Coverage
2.5% 0.898 0.419 0.953
5% 0.900 0.405 0.950
10% 0.899 0.397 0.948
25% 0.899 0.383 0.949
50% 0.898 0.400 0.951
(a) Plug-in regression
Validation Prop. Median 5% 95%
Normalized Bias
2.5% -0.003 -0.052 0.057
5% -0.002 -0.064 0.058
10% -0.001 -0.047 0.053
25% -0.004 -0.052 0.053
50% -0.003 -0.057 0.052
Coverage
2.5% 0.905 0.858 0.966
5% 0.934 0.908 0.958
10% 0.943 0.928 0.956
25% 0.946 0.935 0.957
50% 0.948 0.936 0.959
(b) Debiased regression
Table A13: Summary statistics for normalized bias and coverage using policy topic as a covariate for GPT-5-nano, varying the size of the validation sample.
Figure A12: Cumulative distribution function of mean square error for the bias-corrected estimator against validation-sample only estimator using policy topic as the validation sample size varies.